An unknown amount and concentration of KOH was added to a buffer solution with pH 6.32. We know the amount of conjugate base has tripled relative to the amount of weak acid. What is the new pH of the buffer?
1
Expert's answer
2018-01-04T03:36:51-0500
V = n · Vpart
V (H2O) = 4 · 1.9 x 10^-5 = 7.6 x 10^-5 m3
V (C2H5OH) = 1 · 5.4 x 10^-5 = 5.4 x 10^-5 m3
Vtotal = 7.6 x 10^-5 + 5.4 x 10^-5 = 0.00013 m3